# Weylian.bib

### v2

@@ 1 Introduction
Let $M$ be a simply connected, orientable smooth manifold. A Riemannian metric $g$ on $M$ is called a metric. Its associated conformal metric on $M$ is
@@
$\{F_{1},F_{2}\}=\sum_{i=1}^{n}\left(\frac{\partial F_{1}}{\partial p_{i}}\frac{\partial F_{2}}{\partial x^{i}}-\frac{\partial F_{1}}{\partial x^{i}}\frac{\partial F_{2}}{\partial p_{i}}\right)\,.$
@@
is the naturally associated multi-Hamiltonian. We call two multi-Hamiltonians $\mathcal{H}$ and $\mathcal{H}^{\prime}$ conformally related, if there is a function $\Omega\in\mathcal{C}^{\infty}(M)$ such that
@@
We say that a multi-Hamiltonian is non-degenerate if $\dim(\mathcal{V})=n+2$, c.f. [KMK2007, KKM-1, KKM-2, KKM-3] for instance.
Non-degenerate multi-Hamiltonians that are superintegrable, i.e. that admit a certain, large number of independent functions $T^{*}M\to\mathbb{R}$ that Poisson commute with the (full) multi-Hamiltonian, have received considerable attention in the literature, e.g. [KKM2018, Kress&Schoebel, Kress07, Capel_phdthesis, KSV2023, PKM2013]. More precisely, consider non-degenerate, second-order superintegrable systems, which are introduced properly below. These systems are classified in dimensions two and three [Evans1990, KKPM2001, KKM-1, KKM-2, Capel&Kress, Kress&Schoebel]. Cross-relations have been found to so-called quadratic algebras, and İnönü-Wigner algebra contractions have been used to construct examples, e.g. [KM2014, Capel&Kress&Post]. Non-degenerate second-order superintegrability has also been related to hypergeometric polynomials organised in the Askey-Wilson scheme [PKM2013, PKM2011].
@@
In recent years, considerable advances have been made in the understanding of the geometric structures underlying maximally superintegrable systems. For instance, for two-dimensional non-degenerate systems in the Euclidean plane, [Kress&Schoebel] establishes an algebraic geometric classification, associating a unique planar line triple arrangement to every superintegrable system. For three-dimensional systems, [Capel&Kress, Capel_phdthesis] provides an explicit description of the conformal classes of these systems using algebraic varieties formed from representations of the conformal group. To facilitate an extension of the classification to higher dimensions, a geometric characterisation of non-degenerate systems has recently been put forward, cf. [KSV2023, KSV2024, KSV2024_bauhaus], proposing a tensorial formalism that makes higher dimensions accessible to an efficient investigation. It revealed that the Riemannian metrics underlying such superintegrable systems are often Hessian metrics [AV2025, CV2025] and that examples on constant curvature spaces arise from solutions to the Witten-Dijkgraaf-Verlinde-Verlinde (WDVV) equation [KSV2024, Vollmer2024].
Stäckel transformations, also known as coupling constant metamorphosis, are a well-known transformation of second-order superintegrable systems, which naturally embeds into the tensorial framework mentioned earlier [KSV2024, KSV2024_bauhaus]. The first of these references argues that these systems should be considered as structures on conformal manifolds (called c-superintegrability therein).
@@
The purpose of the present paper is to clarify the concept of c-superintegrability by shedding light onto it from the angle of Weylian geometry, closing a gap in the literature. We explain how an abundant superintegrable system precisely gives rise to a Weyl structure. We furthermore extend the definition of abundant superintegrable systems to Weylian manifolds, including the case of $2$-dimensional systems. Note that while the ideas of [KSV2024] work in general dimension, there is a focus on $n\geq 3$ regarding the abundant structural equations and c-superintegrability. As we are going to see, c.f. Definitions 4 and 5 below, subtleties arise if the underlying Weylian geometry is fixed in advance. In contrast, the abundant structure itself defines a preferred Weylian structure, which may differ from a given one. We characterise this difference by invariant data. In Section 3.1, we review and improve the invariant structural equations for dimensions $n\geq 3$, c.f. [KSV2024]. In Section 3.1, we determine the analogous invariant equations for the structural equations in dimension $n=2$, based on the structural equations given in [KSV2024_bauhaus], which differ significantly from those in higher dimensions. We conclude the paper in Section 3.3 showing that abundant superintegrable systems can be realised as semi-Weyl structures. These structures have recently received some attention in the context of statistical manifolds and affine hypersurface geometry, e.g. [BN2022, Matsuzoe2001, Matsuzoe2010].
@@
The paper is organised as follows: In the following section, we introduce the main objects of the paper, namely Weylian manifolds and abundant superintegrable systems in subsections 2.1 and 2.2, respectively. In Section 3, the main discussion of the paper will take place. Specifically, in Section 3.1, we discuss the case of systems of dimension $n>2$, while Section 3.2 is devoted to the case of $2$-dimensional systems, for which the underpinning structural equations are considerably different from higher dimensions. Finally, we use so-called semi-Weyl structures (a generalisation of Weyl structures) to encode abundant superintegrable systems, see Section 3.3.
@@ 2.1 Weylian manifolds
The purpose of the present section is to introduce Weylian manifolds, also known as Weyl manifolds, e.g. [Folland1970, Norden1976, CP1999, BFM2023]. Following [MS2020], we use the attribute Weylian in order to avoid confusion with the concept of Weyl manifolds in Cartan geometry, c.f. [CM2023, CS2009]. The ensuing twofold definition follows [MS2020], see also the references therein.
@@
    A Weylian structure $(M,\Phi)$ is a differentiable manifold $M$ together with a Weylian metric $\Phi$ on $M$, i.e. an equivalence class of pairs $(g,\varphi)$ consisting of a (pseudo-)Riemannian metric $g$ and a real-valued differential $1$-form $\varphi$, identified under
@@
This statement goes back to Weyl [Weyl1918, WeylGes]. For more details on these definitions and their equivalence, see [MS2020]. In the remainder of the paper, we are going to speak simply of a Weylian manifold when we consider Weyl(ian) structures and do not intend to specify a particular definition.
@@
Exactness and closedness of a Weylian manifold: A Weyl structure is exact, if the connection $D$ is the Levi-Civita connection of some metric in $c$, and closed, if this property holds in the neighbourhood of each point. Equivalently, these properties hold, if $\theta_{g}$ is exact or closed, respectively, for some (and then any) element of $c$, c.f. [MP2024, BFM2023]. We define exactness and closedness for Weylian structures accordingly.
@@ 2.2 Abundant superintegrable systems
A superintegrable system, or, more precisely speaking, a second-order (maximally) conformally superintegrable system, is a Riemannian manifold $(M,g)$ with a natural Hamiltonian $H:T^{*}M\to\mathbb{R}$ admitting $22n-$ 2 additional functions $F^{(\alpha)}:T^{*}M\to\mathbb{R}$, $21\leq\alpha\leq 2n-$ 2, such that (with $F^{(0)}:=H$):
@@
The first point in the above list is the second-order property, while the number $22n-$ 2 indicates the maximality of the superintegrable system. The condition (3) implies that the functions $F^{(\alpha)}$, $20\leq\alpha\leq 2n-$ 2, are constant when restricted to Hamiltonian trajectories on the zero locus of $H$, i.e. the solution curves $\gamma$ of Hamilton’s system of equations,
@@
Note that the condition (3) is a cubic polynomial in the fibre coordinates $p$ (called momenta) on $T^{*}M$. It is easily checked that it has two homogeneous components (in the momenta), namely a cubic and a linear one. A straightforward computation shows that the cubic homogenous part of the polynomial is equivalent to the condition of a conformal Killing tensor,
@@
$dW=\hat{K}(dV)-V\,\rho-dV\wedge\rho\,,$
where $\hat{K}$ denotes the endomorphism naturally associated to $K$ (via the inverse of the metric $g$), and where $\wedge$ is the usual wedge product.
Convention 1.
@@
We denote by $\Pi_{(ijk)\circ}$ the projector onto the totally symmetric (in $i,j,k$) and trace-free (w.r.t. contraction in any pair of $i,j,k$) part of a tensor field (which may have further indices over which no symmetrisation nor trace-freeness is enforced). The circle $\circ$, in particular, will mean trace-freeness. Consequently, the projector onto the totally tracefree part is denoted by $\Pi_{\circ}$. By $\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$ we mean the space of symmetric and tracefree tensor sections of rank two.

### v3

@@ 1 Introduction
Let $M$ be a simply connected, orientable smooth manifold of dimension $n$. A Riemannian metric $g$ on $M$ is called a metric. Its associated conformal metric on $M$ is
@@
$\{F_{1},F_{2}\}=\sum_{i=1}^{n}\left(\frac{\partial F_{1}}{\partial x^{i}}\frac{\partial F_{2}}{\partial p_{i}}-\frac{\partial F_{1}}{\partial p_{i}}\frac{\partial F_{2}}{\partial x^{i}}\right)\,.$
@@
is the naturally associated multi-Hamiltonian. We say that a multi-Hamiltonian is non-degenerate if $\dim(\mathcal{V})=n+2$, cf. [KMK2007, KKM-1, KKM-2, KKM-3] for instance.
We call two multi-Hamiltonians $\mathcal{H}$ and $\mathcal{H}^{\prime}$ conformally related, if there is a function $\Omega\in\mathcal{C}^{\infty}(M)$ such that
@@
As one quickly verifies, this poses a strong restriction on the function $\Omega$. Indeed, an arbitrary choice of $\Omega$ clearly does not preserve integrals of the motion. This motivates the introduction of so-called conformal integrals, i.e. functions $F:T^{*}M\to\mathbb{R}$ that are only required to remain constant along Hamiltonian trajectories with $H=0$ (details will be given later).
1.1 Second-order superintegrable systems
Non-degenerate multi-Hamiltonians that are superintegrable, i.e. that admit more than $n$ independent functions $T^{*}M\to\mathbb{R}$ that Poisson commute with the (full) multi-Hamiltonian, have received considerable attention in the literature, e.g. [KKM2018, Kress&Schoebel, Kress07, Capel_phdthesis, KSV2023, PKM2013]. Here, we will limit ourselves to maximally superintegrable systems of second order. By this we mean (multi-)Hamiltonians that commute (with respect to the Poisson bracket) with $22n-$ 2 additional integrals of the motion. Let $H\in\mathcal{H}$. For second-order maximal superintegrability, we require functions $F^{(\alpha)}:T^{*}M\to\mathbb{R}$, $21\leq\alpha\leq 2n-$ 2, with $\{H,F^{(\alpha)}\}=0$ such that
$F^{(\alpha)}(x,p)=\sum_{i,j=1}^{n}K^{(\alpha)\,ij}(x)p_{i}p_{j}+W^{(\alpha)}(x),$
for functions $K^{ij}$ and $W$ on $M$, and such that the tuple $(H,F^{(1)},\dots,F^{(2n-2)})$ is functionally independent. We mention the well-known fact that here only the functions $W^{(\alpha)}$ depend on the choice $H\in\mathcal{H}$. Indeed, the coefficients $K^{(\alpha)\,ij}$ are linked to Killing tensor fields, which allows one to formulate the conditions for second-order (maximal) superintegrability geometrically and independent of the choice of the $F^{(\alpha)}$. This will be discussed in more detail in Section 2.2, along with proper definitions of the structures we consider in this paper. In the present section, our focus is to give a brief review of some results in the theory of second-order maximally superintegrable systems.
Historically, the study of superintegrable systems has its roots in Kepler’s work on the two-body problem, i.e. the motion of a planet around a central star. This system is a $3$-dimensional maximally superintegrable system: the angular momenta and the Laplace-Runge-Lenz vector provide $42n-2=$ 4 functionally independent integrals of the motion in addition to the energy. According to [Evans1990], the close relationship of second-order superintegrability and the existence of several systems of separation coordinates (of the Hamilton-Jacobi equation) was first noted by Sommerfeld in [Sommerfeld]. Separation systems indeed played a major role in the understanding of second-order maximal superintegrability. Using this relationship, a classification of $2$-dimensional systems (i.e. second-order maximally superintegrable systems with two degrees of freedom) was achieved for spaces of constant curvature in [Evans1990, KKPM2001], see also [KKM2018, KKM-1, KKM-2]. For $3$-dimensional systems, partial classification results exist, cf. [KKM-3, KKM-4, CK14, Capel&Kress, Capel_phdthesis] and [N1, N2, N3, N4, N5], for instance. Cross-relations have been found to so-called quadratic algebras. İnönü-Wigner algebra contractions have been used to construct examples, e.g. [KM2014, Capel&Kress&Post]. Non-degenerate second-order superintegrability has also been related to hypergeometric orthogonal polynomials organised in the Askey-Wilson scheme [PKM2013, PKM2011].
The concept of non-degeneracy has played a fruitful role in many of these developments. Non-degenerate multi-Hamiltonians appear to have been introduced in [KMP2000]. A classification of non-degenerate second-order maximally superintegrable systems on conformally flat space is to date achieved in dimensions $2$ and $3$, see [Kress07, KKMW2003, KKPM2001, KKM-2, Kress&Schoebel] and [Capel_phdthesis, Capel&Kress], respectively. In the classification of $3$-dimensional systems, an important ingredient is the so-called “$(5\longrightarrow 6)$-theorem” (Theorem 3 in [KKM-3]), which states that $3$-dimensional non-degenerate second-order maximally superintegrable systems admit an additional linearly independent integral of the motion. More generally, one can define the concept of abundantness for non-degenerate superintegrable systems, cf. [KSV2023]: it formalizes the property of a non-degenerate system (which has $12n-$ 1 functionally independent integrals of the motion, including the Hamiltonian) to admit $n(n+1)/2$ linearly independent integrals of the motion. To the best knowledge of the author, all non-degenerate systems known to date are of this particular type.
@@ 1.1 Second-order superintegrable systems
In recent years, considerable advances have been made in the understanding of the geometric structures underlying maximally superintegrable systems. For instance, for two-dimensional non-degenerate systems in the Euclidean plane, [Kress&Schoebel] establishes an algebraic geometric classification, associating a unique planar line triple arrangement to every superintegrable system. For three-dimensional systems, [Capel&Kress, Capel_phdthesis] provide an explicit description of the conformal classes of these systems using algebraic varieties formed from representations of the conformal group. To facilitate an extension of the classification to higher dimensions, a geometric characterisation of non-degenerate systems has recently been put forward, cf. [KSV2023, KSV2024, KSV2024_bauhaus], proposing a tensorial formalism that makes higher dimensions accessible to an efficient investigation. It revealed that the Riemannian metrics underlying non-degenerate superintegrable systems are often Hessian metrics [AV2025, CV2025]. Examples on spaces of constant sectional curvature arise from solutions to the (possibly non-flat) Witten-Dijkgraaf-Verlinde-Verlinde (WDVV) equation [KSV2024, Vollmer2024].
1.2 Stäckel transformations, coupling constant metamorphosis and conformal superintegrability
We consider the classical Maupertuis-Jacobi transformation. Let $H(x,p)=g^{-1}_{x}(p,p)+V(x)$ be a Hamiltonian with potential $V$, and define a new (free) Hamiltonian by
$\tilde{H}(x,p):=\frac{H(x,p)}{V(x)}.$
Then, by a classical result, the Hamiltonian trajectories of $\tilde{H}$ are the geodesics of $H$ after a transformation of the time coordinate to $\tilde{t}$, with $d\tilde{t}=V(x(t),p(t))\,dt$, cf. [Tsiganov2001]. A similar, but extended transformation has been introduced for integrable systems that admit separable coordinates by Boyer, Kalnins and Miller in [BKM1986]. In our context, we may view it (in a simplified manner [^1] The transformation was originally formulated in terms of separation coordinates and Stäckel multipliers. Thus, the original statement considers collections of $n$ mutually commuting integrals of the motion, which are transformed into commuting integrals of the motion.) as follows: consider a pair of Hamiltonians $H_{V}(x,p)=g^{-1}_{x}(p,p)+V(x)$ and $H_{U}(x,p)=g^{-1}_{x}(p,p)+U(x)$ with integrals of the motion $F_{V}(x,p)=A(x,p)+W_{V}(x)$ respectively $F_{U}(x,p)=A(x,p)+W(x)_{U}$ of second order, where $A=A(x,p)$ is homogeneous of quadratic order in the momenta $p$. Note that $A$ is shared by both integrals. Then the new Hamiltonian $\tilde{H}=U^{-1}H_{V}$ is called the Stäckel transform of $H_{V}$ via $U$. It admits the second-order integral $\tilde{F}=F_{V}+(1-W_{U})\tilde{H}$.
A related concept is coupling constant metamorphosis, which was introduced by Hietarinta et al. in [HGDR1984] for integrable Hamiltonians of the form $H_{\mu}(x,p)=\sum_{i,j=1}^{n}g^{ij}(x)p_{i}p_{j}+V(x)+\mu W(x)$ with potentials $V$ and $W$ and a coupling parameter $\mu$. If $F_{\mu}$ is an integral of $H_{\mu}$, then let
$\tilde{H}_{\nu}(x,p)=W(x)^{-1}\left(\sum_{i,j=1}^{n}g^{ij}(x)p_{i}p_{j}+V(x)+\nu\right)$
where $\nu$ is the new coupling parameter. It follows that $\tilde{F}_{\nu}:=F_{\tilde{H}_{\nu}}$ is an integral of $\tilde{H}_{\nu}$. Note that here we replace the coupling parameter $\mu$ by $\nu$, i.e. we swap, conceptually, the roles of the coupling parameter and the energy. To avoid misunderstandings, we mention that Stäckel transformations and coupling constant metamorphosis are in general not the same. However, they coincide in the case of second-order maximally superintegrable systems, which we consider here. For more information on the relationship of the two transformations, we refer the interested reader to [Post10].
Note that Stäckel transformations perform a conformal rescaling of the Hamiltonian, $H\longrightarrow\frac{H}{\Omega^{2}}$, where $\Omega^{2}$ however has to be a very specific choice of function. Indeed, rescalings of the above type obviously do not, in general, preserve superintegrability of a given Hamiltonian. A very natural question thus is whether a broader framework exists in which we may carry out rescalings via a generic function $\Omega$. The answer is yes: generic conformal rescalings do preserve the so-called second-order (maximal) conformal superintegrability, in which integrals of the motion are replaced by conformal integrals, i.e. functions that are constant along Hamiltonian trajectories with $H=0$. (Note that $H$ is an integral of the motion and hence $H=0$ is preserved along trajectories.) A precise definition of conformally superintegrable systems is given in Section 2.2. Conformal superintegrability aligns very naturally with the (aforementioned) tensorial framework developed in [KSV2024, KSV2024_bauhaus]. Indeed, the first of these references argues that these systems should be considered as structures on conformal manifolds (called c-superintegrability therein).
@@ 1.2 Stäckel transformations, coupling constant metamorphosis and conformal superintegrability
The purpose of the present paper is to clarify the concept of c-superintegrability by shedding light onto it from the angle of Weylian geometry, closing a gap in the literature. We explain how an abundant superintegrable system gives rise to a Weyl structure. We furthermore extend the definition of abundant superintegrable systems to Weylian manifolds, including the case of $2$-dimensional systems. (Note that, while the ideas of [KSV2024] work in general dimension, there is a focus on $n\geq 3$ regarding the abundant structural equations and c-superintegrability.) As we are going to see, cf. Definitions 6 and 7 below, subtleties arise if the underlying Weylian geometry is fixed in advance. In contrast, the abundant structure itself defines a preferred Weylian structure on $(M,c)$, which may differ from a given one. We characterise this difference by invariant data. In Section 3.1, we review and improve the invariant structural equations for dimensions $n\geq 3$, cf. [KSV2024]. In Section 3.1, we determine the analogous invariant equations for the structural equations in dimension $n=2$, based on the structural equations given in [KSV2024_bauhaus], which differ significantly from those in higher dimensions. We conclude the paper in Section 3.3 showing that abundant superintegrable systems can be realised as semi-Weyl structures. These structures have received some attention in the context of statistical manifolds and affine hypersurface geometry, e.g. [BN2022, Matsuzoe2001, Matsuzoe2010], for instance.
@@
The paper is organised as follows: In the next section, we introduce the main objects of the paper, namely Weylian manifolds and abundant superintegrable systems in Sections 2.1 and 2.2, respectively. In Section 3, the main discussion of the paper will take place. Specifically, in Section 3.1, we discuss the case of systems of dimension $n>2$, while Section 3.2 is devoted to the case of $2$-dimensional systems, for which the underpinning structural equations are considerably different from higher dimensions. Finally, we use so-called semi-Weyl structures (a generalisation of Weyl structures) to encode abundant superintegrable systems, see Section 3.3.
@@ 2 Preliminaries
In this section, we introduce the main structures used in this paper, namely Weylian manifolds (Section 2.1) and abundant non-degenerate superintegrable systems (Section 2.2), which geometrically can be formalized as abundant structures (Section 2.2.1).
@@ 2.1 Weylian manifolds
The purpose of the present section is to introduce Weylian manifolds, also known as Weyl manifolds, e.g. [Folland1970, Norden1976, CP1999, BFM2023]. Following [MS2020], we use the attribute Weylian in order to avoid confusion with the concept of Weyl manifolds in Cartan geometry, cf. [CM2023, CS2009]. The ensuing twofold definition follows [MS2020], see also the references therein.
@@
    A Weylian structure $(M,\Phi)$ is a differentiable manifold $M$ together with a Weylian metric $\Phi$ on $M$, i.e. an equivalence class of pairs $(g,\varphi)$ consisting of a Riemannian metric $g$ and a real-valued differential $1$-form $\varphi$, identified under
@@
This statement goes back to Weyl [Weyl1918, WeylGes]. We then find ($X,Y,Z\in\mathfrak{X}(M)$)
$\nabla_{Z}^{(g,\varphi)}(\Omega^{2}g)(X,Y)$ $=2\Omega^{2}\,(d\ln|\Omega|)(Z)\,g(X,Y)-2\varphi(Z)\Omega^{2}\,g(X,Y)$
$=-2\,(\varphi(Z)-(d\ln|\Omega|)(Z))\,g(X,Y),$
i.e. we can set $D=\nabla^{(g,\varphi)}$, yielding a Weyl structure $(M,[g],D)$. However, note that the Weyl structure is not unique: in fact, for any $1$-form $\alpha\in\Omega^{1}(M)$, $(M,[g],D^{\prime})$ with $D^{\prime}_{X}Y:=D_{X}Y+\alpha(X)Y$ is also a Weyl structure. On the level of Weylian structures, this corresponds to a shift of $\varphi$ by $\alpha$, and therefore we may identify Weylian structures and Weyl structures. For more details, see [MS2020]. In the remainder of the paper, we speak simply of a Weylian manifold whenever we do not wish to specify whether we consider a given setup as in Definition 1 as a Weyl or a Weylian structure.
@@
Exactness and closedness of a Weylian manifold: A Weyl structure is exact, if the connection $D$ is the Levi-Civita connection of some metric in $c$, and closed, if this property holds in the neighbourhood of each point. Equivalently, these properties hold, if $\theta_{g}$ is exact or closed, respectively, for some (and then any) element of $c$, cf. [MP2024, BFM2023]. We define exactness and closedness for Weylian structures accordingly.
@@ 2.2 Abundant superintegrable systems
In Section 1, we have mentioned the role of conformal transformations (presented as Stäckel transformations or coupling constant metamorphosis) in the context of second-order superintegrability. We have also indicated that generic conformal rescalings of the Hamiltonian, while not preserving superintegrability, do preserve conformal superintegrability. We also recall that Weyl manifolds allow for conformal changes of the metric. With this in mind, it is obvious that the natural setting for the study of the relationship of Weylian geometry and second-order superintegrability is that of second-order conformally superintegrable systems.
Convention 1.
Note that in the following, in order to keep the discussion concise, we will usually drop the attributes ‘maximal’, ‘second-order’ and ‘conformal’ when discussing second-order maximally conformally superintegrable systems, as there is no risk of confusion.
A superintegrable system, or, more precisely speaking, a second-order maximally conformally superintegrable system, is a Riemannian manifold $(M,g)$ with a natural Hamiltonian $H:T^{*}M\to\mathbb{R}$ admitting $22n-$ 2 additional functions $F^{(\alpha)}:T^{*}M\to\mathbb{R}$, $21\leq\alpha\leq 2n-$ 2, such that (with $F^{(0)}:=H$)
@@
The first point in the above list is the second-order property, and the number $22n-$ 2 indicates the maximality of the superintegrable system. Equation (3) implies that the functions $F^{(\alpha)}$, $20\leq\alpha\leq 2n-$ 2, are constant when restricted to Hamiltonian trajectories on the zero locus of $H$, i.e. the solution curves $\gamma$ of Hamilton’s system of equations,
@@
Consider (3) and recall that we use canonical Darboux coordinates $(x,p)$. As indicated, for each $\alpha$, $\varrho^{(\alpha)}$ is a function on $T^{*}M$ which is polynomial in the momenta $p$. It is easily checked that the left hand side of (3) is a cubic polynomial in momenta (with coefficients that depend on the position $x$). In fact, it has two homogeneous components (in the momenta), namely a cubic and a linear one. The same has to be true for the right hand side. Therefore, it follows that $\varrho$ has to be linear in $p$, and hence there is a $1$-form $\rho\in\Omega^{1}(M)$ with $\varrho=\rho(p)$. A straightforward computation shows that the cubic homogenous (in momenta) part of the polynomial condition (3) is equivalent to the condition of a conformal Killing tensor,
@@
$dW=\hat{K}(dV)+V\,\rho\,,$ (5)
where $\hat{K}$ denotes the endomorphism naturally associated to $K$ (via the inverse of the metric $g$).
Consider a conformal integral, i.e. a function $F:T^{*}M\to\mathbb{R}$ such that $\{H,F\}=\varrho\,H$ as above. Denote the conformal Killing tensor associated to $F$ by $K$. It follows that also $F^{\prime}=F-\frac{1}{n}\,\mathrm{tr}_{g}(K)\,H$ is a conformal integral,
$\{H,F^{\prime}\}=\{H,F\}-\frac{1}{n}\,\{H,\mathrm{tr}_{g}(K)\}\,H=\left(\varrho-\frac{1}{n}\,(d\mathrm{tr}_{g}(K))(p)\right)\,H=:\varrho^{\prime}\,H\,.$
As a result, the quadratic terms of $F^{\prime}$ correspond to a trace-free conformal Killing tensor. This allows us to agree to the following convention.
Convention 2.
@@
We denote by $\Pi_{(ijk)\circ}$ the projector onto the totally symmetric (in $i,j,k$) and trace-free (w.r.t. contraction in any pair of $i,j,k$) part of a tensor field (which may have further indices over which no symmetrisation nor trace-freeness is enforced). The circle $\circ$, in particular, will mean trace-freeness. Consequently, the projector onto the totally tracefree part is denoted by $\Pi_{\circ}$. By $\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$ we denote the space of sections in the symmetric, trace-free tensors of rank two.
We return to (5). Our aim is now to explain why the concept of a second-order maximally superintegrable system is well-defined for a multi-Hamiltonian system $\mathcal{H}$. To this end, recall that in the integrals $F$, of the form $F=\sum_{i,j=1}^{n}K^{ij}p_{i}p_{j}+W$, only the scalar part $W$ depends on the choice $H\in\mathcal{H}$. Indeed, the condition of a conformal Killing tensor only relies on the metric $g$. Hence, having fixed $H\in\mathcal{H}$, we may integrate (5) for $W$, provided that the integrability condition for (5) holds. It is obtained by applying the differential (exterior derivative) to (5). We find the conformal version of the so-called Bertrand-Darboux condition [bertrand_1857, darboux_1901]
$d(\hat{K}(dV))+V\,d\rho+dV\wedge\rho=0\,.$ (6)
In other words, once we have specified an element $H\in\mathcal{H}$ and a trace-free conformal Killing tensor such that (6) holds, we may integrate for $W$. We introduce the space
$\mathcal{K}:=\{K\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\,|\,$ $K$ is a trace-free conformal Killing tensor
[…]
Given an element of $\mathcal{K}$, we are hence able to recover the integrals for the multi-Hamiltonian $\mathcal{H}$, i.e. we obtain the integrals for any choice of $H\in\mathcal{H}$. This ensures that the concept of second-order (maximal conformal) superintegrability is well-defined for a multi-Hamiltonian. If the multi-Hamiltonian is non-degenerate, we say that the pair $(\mathcal{H},\mathcal{K})$ is a non-degenerate (second-order maximally conformally) superintegrable system.
Non-degenerate superintegrable systems have been studied intensively in the literature, cf. [KKM2018] for an overview of many results on the subject. In [KSV2023, KSV2024, KSV2024_bauhaus], a geometric approach is pursued that is particularly suited for higher dimensions. Specifically, assume that $\mathcal{K}$ is irreducible, i.e. that the space of associated endomorphisms,
$\mathcal{K}^{\sharp}:=\{A\in\mathrm{End}(TM)\,|\,g(A(X),Y)=K(X,Y)\text{ for some $K\in\mathcal{K}$}\,\}\,,$
is irreducible, in the sense that pointwise on an open and dense subset of $M$ they do not have a non-trivial invariant subspace in common. Then it is shown in [KSV2024] that we may write
[…]
$T_{ijk}=S_{ijk}+\bar{t}_{i}g_{jk}+\bar{t}_{j}g_{ik}+\bar{t}_{k}g_{ij}$
where $S_{ijk}$ and $\bar{t}_{i}$ are components of a totally symmetric and trace-free tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ})$ and of a closed $1$-form $\bar{t}\in\Omega^{1}(M)$, respectively. Locally, $\bar{t}$ is therefore the differential of a function, say $\bar{t}=dt$, which we will draw upon in the following.
We now introduce the concept of abundant superintegrable systems. As mentioned earlier, this concept can be traced back to [KKM-3], from which it follows that all $3$-dimensional non-degenerate superintegrable systems are abundant, i.e. they admit an additional (linearly independent) integral of the motion.
Definition 2.
A non-degenerate (second-order maximally conformally) superintegrable system is said to be abundant if $\mathcal{K}$ has dimension $\frac{1}{2}\,n(n+1)-1=\frac{1}{2}\,(n-1)(n+2)$.
2.2.1 Abundant structures
Abundant superintegrable systems are a convenient model case for the study of non-degenerate systems because all known examples of non-degenerate superintegrable systems are, in fact, abundant. Following [KSV2023, KSV2024, KSV2024_bauhaus] and [CV2025], we will now introduce a geometric structure, called abundant structure, to encapsulates an abundant superintegrable system in a geometric manner. It is based on the structural equations determined in [KSV2024] and [KSV2024_bauhaus] for abundant non-degenerate superintegrable systems in dimensions $n\geq 3$ and, respectively, $n=2$. The relationship of Definition 2 above and Definition 3 below is as follows:
  • for a given abundant non-degenerate superintegrable system, the tensors $S$, $\bar{t}$ and $\tau$ satisfy (locally, where $\bar{t}=dt$ for a function $t$) the conditions in Definition 3.
  • for a given abundant structure, we can solve (by integration) the equations [^2] In the case $n\geq 3$, we add Equation (10) below to the system (8) in order to define $\tau$. in Definition 3 for unique solutions $S$, $\bar{t}$ and $\tau$ in a neighborhood of a given point $x\in M$. We can then integrate (7) for $V$ up to the choice of $n+2$ integration constants. It was proven in [KSV2024] that for a generic choice of such a solution we can then integrate (6) for $K$ (subject to $\frac{1}{2}(n-1)(n+2)$ integration constants), and that, from the resulting space $\mathcal{K}$, we can choose $22n-$ 2 elements $K^{(\alpha)}$ such that $(H,F^{(1)},\dots,F^{(2n-2)})$ is a functionally independent collection of conformal integrals of the motion, where $F^{(\alpha)}=\sum_{i,j=1}^{n}K^{(\alpha)\,ij}p_{i}p_{j}+W^{(\alpha)}$ with $W^{(\alpha)}$ being the corresponding solution of (5).
Equivalently, we may say that locally the concept of abundant structures and of abundant non-degenerate superintegrable systems coincide.